OpenStudy (anonymous):

(5z^4)^2(2z^6)^3 Simplify

OpenStudy (harsimran_hs4):

25z^8 * 8z^18 \[(x ^{a})^{b} = x ^{a+b}\]

OpenStudy (harsimran_hs4):

clear till now?

OpenStudy (campbell_st):

the problem becomes \[5^2 \times (z^4)^2 \times 2^3 \times (z^6)^3\] the index law for power of a power is which you need to do 1st \[(x^a)^b = x^{a \times b}\] then when you have done that the index law for multiplication of the same base is add the powers which you need to do 2nd \[x^a \times x^b = x^{a + b}\]

OpenStudy (anonymous):

so then i would do (z^4)^2 *(Z^6)^2

OpenStudy (harsimran_hs4):

yes this is fine for variable z but do keep track of the constants